# references.bib

### v1

@@ 1 Introduction
Navigation is an important part of robotics and the focus of this research. There are many different types of approaches to learning to navigate, including reinforcement learning (RL). In particular, hierarchical reinforcement learning (HRL) is a form of reinforcement learning that aims to overcome some of the shortcomings of traditional RL in environments with sparse reward schemes by breaking down tasks into sub-tasks and creating a hierarchy. In this research, we study HRL’s ability to solve complex spatial navigation problems by forming sub-tasks. This research contributes to the scientific community by providing a comparison between standard reinforcement learning and hierarchical reinforcement learning, an evaluation of different ways to create sub-goals, a comparison of manual and automatic sub-goal creation, and a study of the frequency of sub-goals on performance.
@@ 5.5 Experiment 4: Performance Evaluation on Frequency of Terminations
One of the expected strengths of applying Hierarchical Reinforcement Learning is its ability to break down tasks into sub-tasks and create sub-goals. In this experiment termination created sub-goals will be used. Sub-goals are an important part of the hierarchy created by the algorithm because they define how the task is divided into parts. Because of it’s importance this experiment tests the effects of the frequency of termination on the performance of the algorithm. The frequency of termination determines how many sub-goals are created, as the sub-goal is the state an option reaches before switching to the next option. Additionally if the termination function does not create enough sub-goals then the problem won’t be split up correctly and one option could end up covering the whole task, also known as degerenating. The termination regulation, $\phi$ value was added to the termination function update equation to control the frequency of termination as shown in Equation 2. It is hypothesized as the $\phi$ shrinks, the algorithm will be able to create better termination sub-goals because the smaller $\phi$ will prevent the options from growing too large and degenerating. To test this hypothesis, the Option Critic will be run on the four-rooms maze while varying the $\phi$. All the hyperparameters will be fixed while the $\phi$ is varied. $\phi$ will be varied between 0 and 0.1 with increments of 0.01. This experiment will be performed on the four-rooms maze. The maximum number of steps used in this experiment is 500000 steps. This experiment was run five times to verify the results, and so a statistical test could be performed. An ANOVA test was used to test the differences between each termination value for each of the metrics: convergence time, option lengths, and path length. The results of this experiment are shown in Table 8. To better understand the results, several graphs of one of the experiments are constructed and shown below in Figures 10, 11, and 12. In Figure 10, the option lengths are shown for each of the termination regulation values. In Figure 11, the path lengths are shown as the algorithm is trained for each of the tested values. In Figure 12, the shortest path is shown for each of the test values.
@@ 6.3 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
The idea behind sub-goals is to break down tasks into smaller tasks, such that the smaller tasks are easier for the policies to solve as they are not as complicated. But it is important that the task is split evenly. If not, one of the tasks will be more difficult and would remove the purpose of creating the smaller tasks to make easier tasks. Since it is important to break down the tasks correctly, it may seem best to set the sub-goals by hand to ensure the task is broken down correctly. To manually set the best possible sub-goals for this experiment, each doorway and the goal were chosen as the sub-goals. These were chosen because they evenly break down the task. This is similar to the originally designed four-rooms maze task, where each room was assigned a different option, with the doorway being the sub-goals [SUTTON]. As seen in Table 7 when the algorithm uses the manually set sub-goals it takes about twice as long to converge and it’s path length is about ten steps longer. These results support the hypothesis stating the algorithm will perform better if the sub-goals are set by the algorithm. These results support the hypothesis suggesting the algorithm is able to break down the tasks adequately, thus hand-set sub-goals are not required and even hinder the process as they impede exploration as most of the exploration in the Option Critic comes from trying different options and correspondingly pick a different action. The consideration is that if the sub-goals are set by hand, the algorithm wouldn’t be considering different options in the same state, explaining why automatic sub-goals produce better results. In summary the automatic sub-goals perform better because they allow the algorithm to explore more not because the sub-goals are necessarily better.
@@ 6.4 Experiment 4: Performance Evaluation on Frequency of Terminations
Theoretically, sub-goals are advantageous because they enable an algorithm to break down tasks into several smaller, simpler sub-tasks, which are easier for the algorithm to solve. These results support this hypothesis because the Option Critic is able to outperform the PPO as shown in Table 3. But Option Critics’ increased performance could come from other factors, in addition to sub-goals. In Table 8, it shows how as the termination regulation value increases from 0.00 to 0.10, the convergence time also increases. In Figure 11, it shows how as the termination regulation value increases, the sub-goals are longer and there are fewer sub-goals, up to the point where there are almost no sub-goals. In general, as the number of sub-goals decreases, the performance decreases, showing how it is important to have sub-goals. This might suggest that it would be best to terminate in every state, but as seen in experiment 3, this can also lead to poor performance.

### v2

@@ 1 Introduction
Navigation is an essential part of robotics. Every robot needs to navigated whether it is an arm in a factory moving from part to part or a roomba navigating the room to clean. But navigation problems can be extremely complex, this research aims to find out how to deal with this complexity. The proposed solution to the navigation problem is reinforcement learning, but it can struggle with tasks that have sparse reward structures. One method to deal with this complexity is to build a hierarchy. This comes from nature as humans build hierarchies to solve problems. For example if someone was going to bake a cake, they would first break it into to parts, gathering the ingredients, mixing the ingredients, etc. This is a form of build a hierarchy by breaking the task down into smaller sub-tasks. This can be used in conjunction with reinforcement learning to better solve complex tasks.
Hierarchical reinforcement learning (HRL) is a form of reinforcement learning that aims to overcome some of the shortcomings of traditional RL in environments with sparse reward schemes by breaking down tasks into sub-tasks and creating a hierarchy. In this research, we study HRL’s ability to solve complex spatial navigation problems by forming sub-tasks. This research contributes to the scientific community by providing a comparison between standard reinforcement learning and hierarchical reinforcement learning, an evaluation of different ways to create sub-goals, a comparison of manual and automatic sub-goal creation, and a study of the frequency of sub-goals on performance.
@@ 3.3 Option Critic Architecture
The Option Critic Architecture itself provides no guarantee of convergence, but it is built off of the Option Architecture, which guarantees the convergence of the option policy if every action in every state is executed infinitely often [SUTTON]. There is one main reason this guarantee does not carry over into the Option Critic Architecture, which is that the Option Architecture does not learn the sub-goals and assumes they are set beforehand. Since the Option Critic Architecture learns the sub-goals, the guarantee of convergence does not apply. Meaning if the sub-goals are not correct, the algorithm may not converge, but in practice, this is rare as the algorithm is fairly robust.
@@ 5.5 Experiment 4: Performance Evaluation on Frequency of Terminations
One of the expected strengths of applying Hierarchical Reinforcement Learning is its ability to break down tasks into sub-tasks and create sub-goals. In this experiment termination created sub-goals will be used. Sub-goals are an important part of the hierarchy created by the algorithm because they define how the task is divided into parts. Because of it’s importance this experiment tests the effects of the frequency of termination on the performance of the algorithm. The frequency of termination determines how many sub-goals are created, as the sub-goal is the state an option reaches before switching to the next option. Additionally if the termination function does not create enough sub-goals then the problem won’t be split up correctly and one option could end up covering the whole task, also known as degerenating. The termination regulation, $\phi$ value was added to the termination function update equation to control the frequency of termination as shown in Equation 2. It is hypothesized as the $\phi$ shrinks, the algorithm will be able to create better termination sub-goals because the smaller $\phi$ will prevent the options from growing too large and degenerating. To test this hypothesis, the Option Critic will be run on the four-rooms maze while varying the $\phi$. All the hyperparameters will be fixed while the $\phi$ is varied. $\phi$ will be varied between 0 and 0.1 with increments of 0.01. These values were chosen to show how the performance changes as the options grow in size. In the original model, the value of 0.01 was used to incentivize creating larger options. Values larger than 0.01 were chosen to test whether incentivizing the creation of larger options is necessary. This experiment will be performed on the four-rooms maze. The maximum number of steps used in this experiment is 500000 steps. This experiment was run five times to verify the results, and so a statistical test could be performed. An ANOVA test was used to test the differences between each termination value for each of the metrics: convergence time, option lengths, and path length. The results of this experiment are shown in Table 8. To better understand the results, several graphs of one of the experiments are constructed and shown below in Figures 10, 11, and 12. In Figure 10, the option lengths are shown for each of the termination regulation values. In Figure 11, the path lengths are shown as the algorithm is trained for each of the tested values. In Figure 12, the shortest path is shown for each of the test values.
@@ 6.3 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
The idea behind sub-goals is to break down tasks into smaller tasks, such that the smaller tasks are easier for the policies to solve as they are not as complicated. But it is important that the task is split evenly. If not, one of the tasks will be more difficult and would remove the purpose of creating the smaller tasks to make easier tasks. Since it is important to break down the tasks correctly, it may seem best to set the sub-goals by hand to ensure the task is broken down correctly. To manually set the best possible sub-goals for this experiment, each doorway and the goal were chosen as the sub-goals. These were chosen because they evenly break down the task. This is similar to the originally designed four-rooms maze task, where each room was assigned a different option, with the doorway being the sub-goals [SUTTON]. As seen in Table 7 when the algorithm uses the manually set sub-goals it takes about twice as long to converge and it’s path length is about ten steps longer. These results support the hypothesis stating the algorithm will perform better if the sub-goals are set by the algorithm. These results support the hypothesis suggesting the algorithm is able to break down the tasks adequately, thus hand-set sub-goals are not required and even hinder the process as they impede exploration as most of the exploration in the Option Critic comes from trying different options and correspondingly pick a different action. The consideration is that if the sub-goals are set by hand, the algorithm wouldn’t be considering different options in the same state, explaining why automatic sub-goals produce better results. In summary the automatic sub-goals perform better because they allow the algorithm to explore more not because the sub-goals are necessarily better. In this research there was no metric to compare the sub-goals to ideal sub-goals but for future research such metric could be used to better understand the differences between sub-goals created by algorithm and sub-goals set by hand.
@@ 6.4 Experiment 4: Performance Evaluation on Frequency of Terminations
Theoretically, sub-goals are advantageous because they enable an algorithm to break down tasks into several smaller, simpler sub-tasks, which are easier for the algorithm to solve. These results support this hypothesis because the Option Critic is able to outperform the PPO as shown in Table 3. But Option Critics’ increased performance could come from other factors, in addition to sub-goals. In Table 8, it shows how as the termination regulation value increases from 0.00 to 0.10, the convergence time also increases. In Figure 11, it shows how as the termination regulation value increases, the sub-goals are longer and there are fewer sub-goals, up to the point where there are almost no sub-goals. In general, as the number of sub-goals decreases, the performance decreases, showing how it is important to have sub-goals. This might suggest that it would be best to terminate in every state, but as seen in experiment 3, this can also lead to poor performance. This shows how important sub-goals are in most situations, but this may not be true in all cases. As there may be simple problems that do not need to be broken up and thus do not require sub-goals. Furthermore, in simple problems, sub-goals may hinder performance by adding unnecessary complexity to the task.